Mark is 12 years older than Denise. In 6 years time Mark will be twice as old as Denise. What are their ages now?
Answer by mareebsiddiqui(7)

Denise's age = x
Mark's age = x + 12
After 6 Years:
Denise's age = 6 + x
Mark's age = 2(6 + x) = 12 + 2x
So:
Before 6 year's Mark's age - Before 6 year's Denise's age = New Mark's age + New Denise's age
Difference of ages before 6 years = Difference of ages after 6 years
(x + 12) - x = (12 + 2x) - (6 + x)
12 = x - 6
x = 12 - 6 = 6
New ages of both are:
Denise's age = 6 + 6 = 12
Mark's age = 12 + 2x = 12 + 2(6) = 12 + 12 = 24
Verification of answer:
Now Mark's age is 12 years older and twice as Denise's age.